OpenStudy (anonymous):

A SQUARE PATIO HAS AN AREA OF 200 SQUARE FEET. HOW LONG IS EACH SIDE OF THE PATIO TO THE NEAREST 0.05

OpenStudy (imstuck):

If the patio is a square and its area is 200 sq ft, then the one side times another side equals 200, right, cuz that's the area formula? And if it is a square, then both the sides are the same length. So if A = s^2, and A = 200, then 200 = s^2 and s = the square root of 200.

OpenStudy (imstuck):

What is the square root of 200 to the nearest hundredth?

OpenStudy (imstuck):